# Statistical-Computational Trade-offs for Recursive Adaptive Partitioning Estimators

### v2

@@ Appendix B Length of Query Paths
$\mathbb{E}\left\{n_{t+1}~{}|~{}n_{t},\mathcal{F}_{t}\right\}=\mathbb{E}\left\{\frac{n_{t+1}+n_{t+1}\circ\iota}{2}~{}\vline~{}n_{t},\mathcal{F}_{t}\right\}.$ (B.1)
Finally, observe that $n_{t+1}$ and $n_{t+1}\circ\iota$ are simply the numbers of samples in the right and left children of $\mathcal{C}_{t}$ when splitting on $x_{k(\mathcal{C}_{t})}$, which gives $n_{t+1}+n_{t+1}\circ\iota=n_{t}$. Plugging this into (B.1) and using the tower property of conditional expectations completes the proof. ∎
Lemma B.2.
@@
$\mathbb{P}\left\{\left|J(\mathbf{x};\mathcal{D}_{n},\Theta)\right|\geq t\right\}\leq\min\left\{\frac{n}{2^{t}},1\right\}.$ (B.2)
@@
$\mathbb{E}\left\{\left|J(\mathbf{x};\mathcal{D}_{n},\Theta)\right|\right\}\leq\log_{2}n+2.$ (B.3)
@@
Using Lemma B.1, the fact that $n_{0}=n$, and the tower property, we get $\mathbb{E}\left\{n_{t}\right\}=\frac{n}{2^{t}}$. Using Markov’s inequality therefore gives
$\mathbb{P}\left\{n_{t}\geq 1\right\}\leq\frac{n}{2^{t}}.$ (B.4)
Next, note that $\left|J(\mathbf{x};\mathcal{D}_{n},\Theta)\right|\geq t$ implies that $n_{t}\geq 1$, which gives (B.2). To obtain (B.3), we compute
$\begin{split}\mathbb{E}\left\{\left|J(\mathbf{x};\mathcal{D}_{n},\Theta)\right|\right\}&=\sum_{t=1}^{\infty}\mathbb{P}\left\{\left|J(\mathbf{x};\mathcal{D}_{n},\Theta)\right|\geq t\right\}\\ &\leq\sum_{t=1}^{\lfloor\log_{2}n\rfloor}1+\sum_{t=\lfloor\log_{2}n\rfloor+1}^{\infty}\frac{n}{2^{t}}\\ &\leq\lfloor\log_{2}n\rfloor+\sum_{t=0}^{\infty}2^{-t}\\ &\leq\log_{2}n+2.\qed\end{split}$ (B.5)
Appendix C Proof of Theorem 5.1
Proof of Theorem 5.1.
The proof is similar to that of Proposition 4.1 (the lower bound for the XOR function), but we need to adapt the ingredients (Lemma 3.2, Lemma 3.3, and Lemma 4.2) to this more complicated setting. These generalizations are provided in the rest of this section. For the first statement, we make use of Lemma C.2 and bound $\delta$ using Lemma C.1. For the second statement, we make use of Lemma C.3 and bound $\delta$ using Lemma C.1, bearing in mind the constraint on $\delta$. ∎
Lemma C.1.
Under Assumption 3.1, let $f_{0}^{*}$ be a function not satisfying MSP, and suppose $\mathcal{A}_{t}$ is greedy. Let $T\subset[d]\backslash S^{*}_{\textsf{MSP}}$ be a traversal for $\mathcal{S}\backslash\mathcal{S}_{\textsf{MSP}}$ of size $s_{\operatorname{T}}$ and let $s_{\textsf{MSP}}\coloneqq\left|S^{*}_{\textsf{MSP}}\right|$. Then for each fixed query point $\mathbf{x}\in\{\pm 1\}^{d}$, support set $S^{*}$, and random seed $\Theta$, the probability of selecting any covariate in the traversal along the query path $\mathcal{Q}(\mathbf{x};\mathcal{D}_{n},\Theta)$ is upper bounded as $\mathbb{P}_{\mathcal{D}_{n}}\left\{J(\mathbf{x};\mathcal{D}_{n},\Theta)\cap T\neq\emptyset\right\}\leq\frac{s_{\operatorname{T}}(\log_{2}n+2)}{d-s_{\textsf{MSP}}-s_{\operatorname{T}}+1}$.
@@
$T=\left\{s_{\textsf{MSP}}+1,s_{\textsf{MSP}}+1,\ldots,s_{\textsf{MSP}}+s_{\operatorname{T}}\right\}.$ (C.1)
@@
$\left|S\cap T^{(a)}\right|\geq\left|S\cap T\right|-1\geq 1.$ (C.2)
@@
$\mathbb{P}\left\{\chi_{S}=1~{}\vline~{}\mathbf{X}_{i,-T^{(a)}\cup S^{*}_{\textsf{MSP}}}=\mathbf{x}_{i,-T^{(a)}\cup S^{*}_{\textsf{MSP}}}\right\}=\frac{1}{2}.$ (C.3)
Combining this with the fact that the covariates are drawn uniformly from {±$\{\pm 1\}^{d-1}$ as well as the symmetry of the splitting criterion, we see that the distribution of $\mathcal{D}_{n}^{(a)}$ is invariant to permutation of the covariate indices in $[d]\backslash\left(T^{(a)}\cup S^{*}_{\textsf{MSP}}\right)$. As such, $J(\mathbf{x};\mathcal{D}_{n}^{(a)},\Theta)\backslash S^{*}_{\textsf{MSP}}$ is the prefix of a uniformly random permutation of $[d]\backslash\left(T^{(a)}\cup S^{*}_{\textsf{MSP}}\right)$.
@@
$P\left\{b\in J(\mathbf{x};\mathcal{D}_{n}^{(a)},\Theta)~{}\vline~{}|J(\mathbf{x};\mathcal{D}_{n}^{(a)},\Theta)|\right\}\leq\frac{|J(\mathbf{x};\mathcal{D}_{n}^{(a)},\Theta)|}{d-s_{\textsf{MSP}}-s_{\operatorname{T}}}.$ (C.4)
Taking a further expectation and using Lemma B.2 gives
$\begin{split}\mathbb{P}\left\{b\in J(\mathbf{x};\mathcal{D}_{n}^{(a)},\Theta)\right\}&=\mathbb{E}\left\{\mathbb{P}\left\{b\in J(\mathbf{x};\mathcal{D}_{n}^{(a)},\Theta)~{}\vline~{}|J(\mathbf{x};\mathcal{D}_{n}^{(a)},\Theta)|\right\}\right\}\\ &\leq\frac{\log_{2}n+2}{d-s_{\textsf{MSP}}-s_{\operatorname{T}}}.\end{split}$ (C.5)
@@
$\bigcup_{a=1}^{s_{\operatorname{T}}}\left\{J(\mathbf{x};\mathcal{D}_{n},\Theta)\neq J(\mathbf{x};\mathcal{D}_{n}^{(a)},\Theta)\right\}\subset\bigcup_{a=1}^{s_{\operatorname{T}}}\left\{s_{\textsf{MSP}}+a\in J(\mathbf{x};\mathcal{D}_{n}^{(a)},\Theta)\right\}.$ (C.6)
@@
$\mathcal{O}(j,\mathcal{C}_{t},\mathcal{D}_{n})=\mathcal{F}\left(\left\{(X_{ij},Y_{i})\right\}_{i\in I(\mathcal{C})}\right)=\mathcal{O}(j,\mathcal{C}_{t},\mathcal{D}_{n}^{(1)}).$ (C.7)
@@
Step 5: Completing the proof. Suppose $J(\mathbf{x};\mathcal{D}_{n},\Theta)\cap T\neq\emptyset$. If coupling does not hold (i.e., the left hand side of (C.6) is true), then by that same equation, we see that $s_{\textsf{MSP}}+a\in J(\mathbf{x};\mathcal{D}_{n}^{(a)},\Theta)$ for some $a=1,2,\ldots,s_{\operatorname{T}}$. On the other hand if coupling does hold, then picking $s_{\textsf{MSP}}+a\in J(\mathbf{x};\mathcal{D}_{n},\Theta)\cap T$, we have $s_{\textsf{MSP}}+a\in J(\mathbf{x};\mathcal{D}_{n},\Theta)=J(\mathbf{x};\mathcal{D}_{n}^{(a)},\Theta)$. In conclusion, we have
$\begin{split}\mathbb{P}_{\mathcal{D}_{n}}\left\{J(\mathbf{x};\mathcal{D}_{n},\Theta)\cap T\neq\emptyset\right\}&\leq\sum_{a=1}^{s_{\operatorname{T}}}\mathbb{P}\left\{s_{\textsf{MSP}}+a\in J(\mathbf{x};\mathcal{D}_{n}^{(a)},\Theta)\right\}\\ &\leq\frac{s_{\operatorname{T}}(\log_{2}n+2)}{d-s_{\textsf{MSP}}-s_{\operatorname{T}}+1}.\qed\end{split}$ (C.8)
Lemma C.2.
@@
$R\left(\hat{f}(-;\mathcal{D}_{n},\Theta),f^{*}\right)\geq\mathbb{E}_{\mathbf{X}\sim\nu}\left\{\left(\hat{f}(\mathbf{X};\mathcal{D}_{n},\Theta)-f^{*}(\mathbf{X})\right)^{2}\mathbb{1}\left\{J(\mathbf{X};\mathcal{D}_{n},\Theta)\cap T=\emptyset\right\}\right\}.$ (C.9)
@@
$\begin{split}&\mathbb{E}_{\mathbf{X}}\left\{\left(\hat{f}(\mathbf{X};\mathcal{D}_{n},\Theta)-f^{*}(\mathbf{X})\right)^{2}~{}\vline~{}\mathbf{X}_{[d]\backslash T}=\mathbf{x}_{[d]\backslash T}\right\}\\ &=\operatorname{Var}\left\{f_{0}^{*}(\mathbf{X})~{}|~{}\mathbf{X}_{[d]\backslash T}=\mathbf{x}_{[d]\backslash T}\right\}+\left(\hat{f}(\mathbf{x};\mathcal{D}_{n},\Theta)-\mathbb{E}\left\{f_{0}^{*}(\mathbf{X})~{}|~{}\mathbf{X}_{[d]\backslash T}=\mathbf{x}_{[d]\backslash T}\right\}\right)^{2}\\ &\geq\operatorname{Var}\left\{f_{0}^{*}(\mathbf{X})~{}|~{}\mathbf{X}_{[d]\backslash T}=\mathbf{x}_{[d]\backslash T}\right\}.\end{split}$ (C.10)
Making the further observation that the function $\mathbb{1}\left\{J(\mathbf{x};\mathcal{D}_{n},\Theta)\cap T=\emptyset\right\}$ is constant with respect to $\mathbf{x}_{T}$ (i.e., is in the $\sigma$-algebra generated by $\mathbf{X}_{[d]\backslash T}$), we can use conditional expectations to rewrite the right hand side of (C.9) as
$\mathbb{E}_{\mathbf{X}\sim\nu}\left\{\mathbb{E}_{\mathbf{X}}\left\{\left(\hat{f}(\mathbf{X};\mathcal{D}_{n},\Theta)-f^{*}(\mathbf{X})\right)^{2}~{}\vline~{}\mathbf{X}_{[d]\backslash T}\right\}\mathbb{1}\left\{J(\mathbf{X};\mathcal{D}_{n},\Theta)\cap T=\emptyset\right\}\right\}.$ (C.11)
Plugging in the bound (C.10), taking a further expectation with respect to $\mathcal{D}_{n}$ and $\Theta$ shows that
$\mathfrak{R}\left(\hat{f},f^{*}_{0},d,n\right)\geq(1-\delta)\mathbb{E}\left\{\operatorname{Var}\left\{f^{*}(\mathbf{X})~{}|~{}\mathbf{X}_{T}\right\}\right\}.$ (C.12)
@@
$\mathbb{E}\left\{\chi_{S}~{}|~{}\mathbf{X}_{[d]\backslash T}=\mathbf{x}_{[d]\backslash T}\right\}=0.$ (C.13)
@@
$\begin{split}\mathbb{E}\left\{\operatorname{Var}\left\{f^{*}(\mathbf{X})~{}|~{}\mathbf{X}_{T}\right\}\right\}&=\operatorname{Var}\left\{f^{*}(\mathbf{X})\right\}-\operatorname{Var}\left\{\mathbb{E}\left\{f^{*}(\mathbf{X})~{}|~{}\mathbf{X}_{[d]\backslash T}\right\}\right\}\\ &=\operatorname{Var}\left\{f^{*}(\mathbf{X})\right\}-\operatorname{Var}\left\{\sum_{S\in\mathcal{S},S\cap T=\emptyset}\alpha_{S}\chi_{S}\right\}\\ &\geq\sum_{S\in\mathcal{S}^{\prime}}\alpha_{S}^{2}.\qed\end{split}$ (C.14)
Lemma C.3.
@@
$\hat{f}(\mathbf{x};\mathcal{D}_{n},\Theta)=\frac{1}{B}\sum_{b=1}^{B}\tilde{f}(\mathbf{x};\mathcal{D}_{n},\theta_{b}),$ (C.15)
@@
$\begin{split}\mathbb{E}_{\mathcal{D}_{n},\Theta}\left\{\frac{1}{B}\sum_{b=1}^{B}\tilde{f}(\mathbf{x};\mathcal{D}_{n},\theta_{b})\right\}&=\mathbb{E}_{\mathcal{D}_{n},\theta_{1}}\left\{\tilde{f}(\mathbf{x};\mathcal{D}_{n},\theta_{1})\right\}\\ &=\underbrace{\mathbb{E}_{\mathcal{D}_{n},\theta_{1}}\left\{\tilde{f}(\mathbf{x};\mathcal{D}_{n},\theta_{1})\mathbb{1}\left\{J(\mathbf{x};\mathcal{D}_{n},\theta_{1})\cap T=\emptyset\right\}\right\}}_{g(\mathbf{x})}\\ &\quad+\underbrace{\mathbb{E}_{\mathcal{D}_{n},\theta_{1}}\left\{\tilde{f}(\mathbf{x};\mathcal{D}_{n},\theta_{1})\mathbb{1}\left\{J(\mathbf{x};\mathcal{D}_{n},\theta_{1})\cap T\neq\emptyset\right\}\right\}}_{h(\mathbf{x})}.\end{split}$ (C.16)
@@
$\begin{split}&\mathbb{E}_{\mathbf{X},\Theta,\mathcal{D}_{n}}\left\{\left(\hat{f}(\mathbf{X};\mathcal{D}_{n},\Theta)-f^{*}(\mathbf{X})\right)^{2}\right\}\\ &\geq\mathbb{E}_{\mathbf{X}}\left\{\left(g(\mathbf{X})+h(\mathbf{X})-f^{*}(\mathbf{X})\right)^{2}\right\}\\ &=\mathbb{E}_{\mathbf{X}}\left\{\mathbb{E}_{\mathbf{X}}\left\{\left(g(\mathbf{X})+h(\mathbf{X})-f^{*}(\mathbf{X})\right)^{2}~{}\vline~{}\mathbf{X}_{[d]\backslash T}\right\}\right\}.\end{split}$ (C.17)
We next notice that $g(\mathbf{x})$ is constant with respect to changes in $\mathbf{x}_{T}$. This means that, similar to (C.10), the conditional expectation within (C.17) can be lower bounded by a variance term, which we expand as follows:
$\begin{split}&\mathbb{E}_{\mathbf{X}}\left\{\left(g(\mathbf{X})+h(\mathbf{X})-f^{*}(\mathbf{X})\right)^{2}~{}\vline~{}\mathbf{X}_{[d]\backslash T}=\mathbf{x}_{[d]\backslash T}\right\}\\ &\geq\operatorname{Var}_{\mathbf{X}}\left\{h(\mathbf{X})-f^{*}(\mathbf{X})~{}|~{}\mathbf{X}_{[d]\backslash T}=\mathbf{x}_{[d]\backslash T}\right\}\\ &\geq\operatorname{Var}\left\{f^{*}(\mathbf{X})~{}|~{}\mathbf{X}_{[d]\backslash T}=\mathbf{x}_{[d]\backslash T}\right\}\\ &\quad\quad-2\operatorname{Var}\left\{f^{*}(\mathbf{X})~{}|~{}\mathbf{X}_{[d]\backslash T}=\mathbf{x}_{[d]\backslash T}\right\}^{1/2}\operatorname{Var}_{\mathbf{X}}\left\{h(\mathbf{X})~{}|~{}\mathbf{X}_{[d]\backslash T}=\mathbf{x}_{[d]\backslash T}\right\}^{1/2}.\end{split}$ (C.18)
@@
$\begin{split}\mathbb{E}_{\mathbf{X}}\left\{\operatorname{Var}_{\mathbf{X}}\left\{h(\mathbf{X})~{}|~{}\mathbf{X}_{[d]\backslash T}\right\}^{1/2}\right\}&\leq\mathbb{E}_{\mathbf{X}}\left\{h(\mathbf{X})^{2}\right\}^{1/2}\\ &\leq M\delta.\end{split}$ (C.19)
Plugging (C.18) into (C.17) and simplifying using (C.19) gives
$\begin{split}\mathfrak{R}\left(\hat{f},f^{*}_{0},d,n\right)&\geq\mathbb{E}\left\{\operatorname{Var}\left\{f^{*}(\mathbf{X})~{}|~{}\mathbf{X}_{[d]\backslash T}\right\}\right\}-2M\delta\mathbb{E}\left\{\operatorname{Var}\left\{f^{*}(\mathbf{X})~{}|~{}\mathbf{X}_{[d]\backslash T}\right\}\right\}^{1/2}\\ &=\sum_{S\in\mathcal{S},S\cap T=\emptyset}\alpha_{S}^{2}-2M\delta\left(\sum_{S\in\mathcal{S},S\cap T=\emptyset}\alpha_{S}^{2}\right)^{1/2}\\ &\geq\sum_{S\in\mathcal{S}^{\prime}}\alpha_{S}^{2}-2M\delta\left(\sum_{S\in\mathcal{S}^{\prime}}\alpha_{S}^{2}\right)^{1/2},\end{split}$ (C.20)
@@
Appendix D Proof of Proposition 6.1
Proof of Proposition 6.1
We first construct a hypothesis set as follows. Using Lemma D.1, there exists a $1/2$-separated set of $[d-s]^{2^{s-1}}$ of size at least $\exp(2^{s-2}\log((d-s)/4)$, which we denote as $\mathcal{M}$. Now for each $\mathbf{z}\in\mathcal{M}$, let $\mathfrak{T}_{\mathbf{z}}$ be the tree structure comprising splits on covariate $d-s+i+1$ for all nodes at depth $i$ for $i=0,1,\ldots,s-2$, and whose splits at level $s-1$ are determined by $\mathbf{z}$. For every tree structure $\mathfrak{T}_{\mathbf{z}}$, define a leaf parameter vector $\boldsymbol{\mu}$ so that for each leaf $\mathcal{L}_{i}$, the leaf mean satisfies
$\mu_{i}=\begin{cases}1&\text{if}~{}\mathcal{L}_{i}~{}\text{is a right child},\\ -1&\text{if}~{}\mathcal{L}_{i}~{}\text{is a left child}.\end{cases}$ (D.1)
@@
$\mathcal{G}_{\delta}\coloneqq\left\{\delta\cdot f_{\mathbf{z}}\colon\mathbf{z}\in\mathcal{M}\right\},$ (D.2)
@@
$\mathbb{E}\left\{\delta\cdot f_{\mathbf{z}}(\mathbf{X})\right\}=\frac{\delta}{2^{s}}\sum_{j=1}^{2^{s}}\mu_{j}=0,$ (D.3)
@@
$\|\delta\cdot f_{\mathbf{z}}\|_{2}^{2}=\delta^{2}.$ (D.4)
@@
$\phi_{i}(\mathbf{x})=\mathbb{1}\left\{\mathbf{x}\in\mathcal{L}_{2i}\right\}-\mathbb{1}\left\{\mathbf{x}\in\mathcal{L}_{2i-1}\right\},$ (D.5)
@@
$\mathbb{E}\left\{\phi_{i}(\mathbf{X})\phi_{i}^{\prime}(\mathbf{X})\right\}=\begin{cases}2^{-s+1}&\text{if}~{}z_{i}=z_{i}^{\prime},\\ 0&\text{otherwise}.\end{cases}$ (D.6)
@@
$\mathbb{E}\left\{f_{\mathbf{z}}(\mathbf{X})f_{\mathbf{z}^{\prime}}(\mathbf{X})\right\}=2^{-s+1}\sum_{i=1}^{2^{s-1}}\mathbb{1}\left\{z_{i}=z_{i}^{\prime}\right\}=1-d_{H}(\mathbf{z},\mathbf{z}^{\prime}),$ (D.7)
@@
$\left\|\delta\cdot f_{\mathbf{z}}-\delta\cdot f_{\mathbf{z}^{\prime}}\right\|_{2}^{2}=2\delta^{2}d_{H}(\mathbf{z},\mathbf{z}^{\prime})\geq\delta^{2}.$ (D.8)

### v3

@@ Appendix C Length of Query Paths
$\mathbb{E}\left\{n_{t+1}~|~n_{t},\mathcal{F}_{t}\right\}=\mathbb{E}\left\{\frac{n_{t+1}+n_{t+1}\circ\iota}{2}~\vline~n_{t},\mathcal{F}_{t}\right\}.$ (C.1)
Finally, observe that $n_{t+1}$ and $n_{t+1}\circ\iota$ are simply the numbers of samples in the right and left children of $\mathcal{C}_{t}$ when splitting on $x_{k(\mathcal{C}_{t})}$, which gives $n_{t+1}+n_{t+1}\circ\iota=n_{t}$. Plugging this into (C.1) and using the tower property of conditional expectations completes the proof. ∎
Lemma C.2.
@@
$\mathbb{P}\left\{\left|J(\mathbf{x};\mathcal{D}_{n},\Theta)\right|\geq t\right\}\leq\min\left\{\frac{n}{2^{t}},1\right\}.$ (C.2)
@@
$\mathbb{E}\left\{\left|J(\mathbf{x};\mathcal{D}_{n},\Theta)\right|\right\}\leq\log_{2}n+2.$ (C.3)
@@
Using Lemma C.1, the fact that $n_{0}=n$, and the tower property, we get $\mathbb{E}\left\{n_{t}\right\}=\frac{n}{2^{t}}$. Using Markov’s inequality therefore gives
$\mathbb{P}\left\{n_{t}\geq 1\right\}\leq\frac{n}{2^{t}}.$ (C.4)
Next, note that $\left|J(\mathbf{x};\mathcal{D}_{n},\Theta)\right|\geq t$ implies that $n_{t}\geq 1$, which gives (C.2). To obtain (C.3), we compute
$\begin{split}\mathbb{E}\left\{\left|J(\mathbf{x};\mathcal{D}_{n},\Theta)\right|\right\}&=\sum_{t=1}^{\infty}\mathbb{P}\left\{\left|J(\mathbf{x};\mathcal{D}_{n},\Theta)\right|\geq t\right\}\\ &\leq\sum_{t=1}^{\lfloor\log_{2}n\rfloor}1+\sum_{t=\lfloor\log_{2}n\rfloor+1}^{\infty}\frac{n}{2^{t}}\\ &\leq\lfloor\log_{2}n\rfloor+\sum_{t=0}^{\infty}2^{-t}\\ &\leq\log_{2}n+2.\qed\end{split}$ (C.5)
Appendix D Proof of Theorem 5.1
The proof is similar to that of Proposition 4.1 (the lower bound for the XOR function), but we need to adapt the ingredients (Lemma 3.2, Lemma 3.3, and Lemma 4.2) to this more complicated setting. These generalizations are provided in the rest of this section. For the first statement, we make use of Lemma D.2 and bound $\delta$ using Lemma D.1. For the second statement, we make use of Lemma D.3 and bound $\delta$ using Lemma D.1, bearing in mind the constraint on $\delta$. ∎
Lemma D.1.
[…]
@@
$T=\left\{s_{\textsf{MSP}}+1,s_{\textsf{MSP}}+1,\ldots,s_{\textsf{MSP}}+s_{\operatorname{T}}\right\}.$ (D.1)
@@
$\left|S\cap T^{(a)}\right|\geq\left|S\cap T\right|-1\geq 1.$ (D.2)
@@
$\mathbb{P}\left\{\chi_{S}=1~\vline~\mathbf{X}_{i,-T^{(a)}\cup S^{*}_{\textsf{MSP}}}=\mathbf{x}_{i,-T^{(a)}\cup S^{*}_{\textsf{MSP}}}\right\}=\frac{1}{2}.$ (D.3)
Combining this with the fact that the covariates are drawn uniformly from $\{\pm 1\}^{d-1}$ as well as the symmetry of the splitting criterion, we see that the distribution of $\mathcal{D}_{n}^{(a)}$ is invariant to permutation of the covariate indices in $d]\backslash\left(T^{(a)}\cup S^{*}_{\textsf{MSP}}\right)$. As such, $J(\mathbf{x};\mathcal{D}_{n}^{(a)},\Theta)\backslash S^{*}_{\textsf{MSP}}$ is the prefix of a uniformly random permutation of $d]\backslash\left(T^{(a)}\cup S^{*}_{\textsf{MSP}}\right)$.
@@
$P\left\{b\in J(\mathbf{x};\mathcal{D}_{n}^{(a)},\Theta)~\vline~|J(\mathbf{x};\mathcal{D}_{n}^{(a)},\Theta)|\right\}\leq\frac{|J(\mathbf{x};\mathcal{D}_{n}^{(a)},\Theta)|}{d-s_{\textsf{MSP}}-s_{\operatorname{T}}}.$ (D.4)
Taking a further expectation and using Lemma C.2 gives
$\begin{split}\mathbb{P}\left\{b\in J(\mathbf{x};\mathcal{D}_{n}^{(a)},\Theta)\right\}&=\mathbb{E}\left\{\mathbb{P}\left\{b\in J(\mathbf{x};\mathcal{D}_{n}^{(a)},\Theta)~\vline~|J(\mathbf{x};\mathcal{D}_{n}^{(a)},\Theta)|\right\}\right\}\\ &\leq\frac{\log_{2}n+2}{d-s_{\textsf{MSP}}-s_{\operatorname{T}}}.\end{split}$ (D.5)
@@
$\bigcup_{a=1}^{s_{\operatorname{T}}}\left\{J(\mathbf{x};\mathcal{D}_{n},\Theta)\neq J(\mathbf{x};\mathcal{D}_{n}^{(a)},\Theta)\right\}\subset\bigcup_{a=1}^{s_{\operatorname{T}}}\left\{s_{\textsf{MSP}}+a\in J(\mathbf{x};\mathcal{D}_{n}^{(a)},\Theta)\right\}.$ (D.6)
@@
$\mathcal{O}(j,\mathcal{C}_{t},\mathcal{D}_{n})=\mathcal{F}\left(\left\{(X_{ij},Y_{i})\right\}_{i\in I(\mathcal{C})}\right)=\mathcal{O}(j,\mathcal{C}_{t},\mathcal{D}_{n}^{(1)}).$ (D.7)
@@
Step 5: Completing the proof. Suppose $J(\mathbf{x};\mathcal{D}_{n},\Theta)\cap T\neq\emptyset$. If coupling does not hold (i.e., the left hand side of (D.6) is true), then by that same equation, we see that $s_{\textsf{MSP}}+a\in J(\mathbf{x};\mathcal{D}_{n}^{(a)},\Theta)$ for some $a=1,2,\ldots,s_{\operatorname{T}}$. On the other hand if coupling does hold, then picking $s_{\textsf{MSP}}+a\in J(\mathbf{x};\mathcal{D}_{n},\Theta)\cap T$, we have $s_{\textsf{MSP}}+a\in J(\mathbf{x};\mathcal{D}_{n},\Theta)=J(\mathbf{x};\mathcal{D}_{n}^{(a)},\Theta)$. In conclusion, we have
$\begin{split}\mathbb{P}_{\mathcal{D}_{n}}\left\{J(\mathbf{x};\mathcal{D}_{n},\Theta)\cap T\neq\emptyset\right\}&\leq\sum_{a=1}^{s_{\operatorname{T}}}\mathbb{P}\left\{s_{\textsf{MSP}}+a\in J(\mathbf{x};\mathcal{D}_{n}^{(a)},\Theta)\right\}\\ &\leq\frac{s_{\operatorname{T}}(\log_{2}n+2)}{d-s_{\textsf{MSP}}-s_{\operatorname{T}}+1}.\qed\end{split}$ (D.8)
Lemma D.2.
@@
$R\left(\hat{f}(-;\mathcal{D}_{n},\Theta),f^{*}\right)\geq\mathbb{E}_{\mathbf{X}\sim\nu}\left\{\left(\hat{f}(\mathbf{X};\mathcal{D}_{n},\Theta)-f^{*}(\mathbf{X})\right)^{2}\mathbb{1}\left\{J(\mathbf{X};\mathcal{D}_{n},\Theta)\cap T=\emptyset\right\}\right\}.$ (D.9)
@@
[…]
Plugging in the bound (D.10), taking a further expectation with respect to $\mathcal{D}_{n}$ and $\Theta$ shows that
$\mathfrak{R}\left(\hat{f},f^{*}_{0},d,n\right)\geq(1-\delta)\mathbb{E}\left\{\operatorname{Var}\left\{f^{*}(\mathbf{X})~|~\mathbf{X}_{T}\right\}\right\}.$ (D.12)
@@
$\mathbb{E}\left\{\chi_{S}~|~\mathbf{X}_{[d]\backslash T}=\mathbf{x}_{[d]\backslash T}\right\}=0.$ (D.13)
@@
[…]
Lemma D.3.
@@
$\hat{f}(\mathbf{x};\mathcal{D}_{n},\Theta)=\frac{1}{B}\sum_{b=1}^{B}\tilde{f}(\mathbf{x};\mathcal{D}_{n},\theta_{b}),$ (D.15)
@@
$\begin{split}\mathbb{E}_{\mathcal{D}_{n},\Theta}\left\{\frac{1}{B}\sum_{b=1}^{B}\tilde{f}(\mathbf{x};\mathcal{D}_{n},\theta_{b})\right\}&=\mathbb{E}_{\mathcal{D}_{n},\theta_{1}}\left\{\tilde{f}(\mathbf{x};\mathcal{D}_{n},\theta_{1})\right\}\\ &=\underbrace{\mathbb{E}_{\mathcal{D}_{n},\theta_{1}}\left\{\tilde{f}(\mathbf{x};\mathcal{D}_{n},\theta_{1})\mathbb{1}\left\{J(\mathbf{x};\mathcal{D}_{n},\theta_{1})\cap T=\emptyset\right\}\right\}}_{g(\mathbf{x})}\\ &\quad+\underbrace{\mathbb{E}_{\mathcal{D}_{n},\theta_{1}}\left\{\tilde{f}(\mathbf{x};\mathcal{D}_{n},\theta_{1})\mathbb{1}\left\{J(\mathbf{x};\mathcal{D}_{n},\theta_{1})\cap T\neq\emptyset\right\}\right\}}_{h(\mathbf{x})}.\end{split}$ (D.16)
@@
[…]
We next notice that $g(\mathbf{x})$ is constant with respect to changes in $\mathbf{x}_{T}$. This means that, similar to (D.10), the conditional expectation within (D.17) can be lower bounded by a variance term, which we expand as follows:
[…]
@@
[…]
Plugging (D.18) into (D.17) and simplifying using (D.19) gives
[…]
@@
Appendix E Proof of Proposition 6.1
We first construct a hypothesis set as follows. Using Lemma E.1, there exists a $1/2$-separated set of $d-s]^{2^{s-1}}$ of size at least $\exp(2^{s-2}\log((d-s)/4)$, which we denote as $\mathcal{M}$. Now for each $\mathbf{z}\in\mathcal{M}$, let $\mathfrak{T}_{\mathbf{z}}$ be the tree structure comprising splits on covariate ds+i+1d-s+i+1 for all nodes at depth $i$ for $i=0,1,\ldots,s-2$, and whose splits at level s1s-1 are determined by $\mathbf{z}$. For every tree structure $\mathfrak{T}_{\mathbf{z}}$, define a leaf parameter vector $\boldsymbol{\mu}$ so that for each leaf $\mathcal{L}_{i}$, the leaf mean satisfies
$\mu_{i}=\begin{cases}1&\text{if}~\mathcal{L}_{i}~\text{is a right child},\\ -1&\text{if}~\mathcal{L}_{i}~\text{is a left child}.\end{cases}$ (E.1)
@@
$\mathcal{G}_{\delta}\coloneqq\left\{\delta\cdot f_{\mathbf{z}}\colon\mathbf{z}\in\mathcal{M}\right\},$ (E.2)
@@
$\mathbb{E}\left\{\delta\cdot f_{\mathbf{z}}(\mathbf{X})\right\}=\frac{\delta}{2^{s}}\sum_{j=1}^{2^{s}}\mu_{j}=0,$ (E.3)
@@
$\|\delta\cdot f_{\mathbf{z}}\|_{2}^{2}=\delta^{2}.$ (E.4)
@@
$\phi_{i}(\mathbf{x})=\mathbb{1}\left\{\mathbf{x}\in\mathcal{L}_{2i}\right\}-\mathbb{1}\left\{\mathbf{x}\in\mathcal{L}_{2i-1}\right\},$ (E.5)
@@
$\mathbb{E}\left\{\phi_{i}(\mathbf{X})\phi_{i}^{\prime}(\mathbf{X})\right\}=\begin{cases}2^{-s+1}&\text{if}~z_{i}=z_{i}^{\prime},\\ 0&\text{otherwise}.\end{cases}$ (E.6)
@@
$\mathbb{E}\left\{f_{\mathbf{z}}(\mathbf{X})f_{\mathbf{z}^{\prime}}(\mathbf{X})\right\}=2^{-s+1}\sum_{i=1}^{2^{s-1}}\mathbb{1}\left\{z_{i}=z_{i}^{\prime}\right\}=1-d_{H}(\mathbf{z},\mathbf{z}^{\prime}),$ (E.7)
@@
$\left\|\delta\cdot f_{\mathbf{z}}-\delta\cdot f_{\mathbf{z}^{\prime}}\right\|_{2}^{2}=2\delta^{2}d_{H}(\mathbf{z},\mathbf{z}^{\prime})\geq\delta^{2}.$ (E.8)